order: Aquamephyton 15 mg IM stat

Supply: Aquamephyton 10 mg per mL

Give: _________ mL(s) (Round to nearest whole number)
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required.

       Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

Here

     Desired dose(D) = 15 mg.

     Dose in hand(H) = 10 mg.

        Quantity(Q) = 1 ml

As per above formula

    Volume(ml) of drug required = D/H * Q

                                                  = 15/10 * 1

                                                  = 1.5 * 1

                                                  = 1.5 ml.